Please help me to solve this,
For what values of K will the lines 2x-3y=5 and 4x=9+Ky be perpendicular?

Answer by stanbon(75887)

Please help me to solve this,
For what values of K will the lines 2x-3y=5 and 4x=9+Ky be perpendicular?
------------------
Put each equation in slope-intercept form:
y = (2/3)x - (5/3)
slope = 2/3
-------------------
y = (4/K)y - (9/K)
slope = 4/K
-----
Note:: For the lines to be perpendicular the product of the
slopes must be -1.
---
(2/3)(4/K) = -1
8/(3K) = -1
3K = -8
K = -8/3
-------------
Cheers,
Stan H.
-------------------